OpenStudy (anonymous):

the diagonal of a rectangle measures 45 cm. the width of the rectangle is 27 cm. find the length of the rectangle.

OpenStudy (anonymous):

you have a right angle triangle here with hypotenuse=45cm. thus \[c^2=a^2+b^2\]\[l^2=(45)^2-(27)^2=1296\]\[l=36\]